OpenStudy (starryknight):

can someone please explain how to do this ...please explain in steps Find the sum of the first 26 terms of the following sequence: -12, -13, -14 . . . -1274 -637 -98 -312

OpenStudy (agent0smith):

You can use an arithmetic sequence formula to find the sum. First term is -12, common difference is -1, number of terms is 26.

OpenStudy (nirmalnema):

s(n)=n/2 (2a+(n-1)d) n=no of terms.=26 a=first term=-12 d=common difference=-13-(-12)=-1 plug in all the values 26/2(-24+(26-1)(-1)) =13(-49)=-637

OpenStudy (starryknight):

thank you sooo much :)

OpenStudy (nirmalnema):

its ok .. ur welcome.!! :)

OpenStudy (agent0smith):

lol. You could let her at least try to do some of the work...

OpenStudy (starryknight):

thank you both